OpenStudy (anonymous):

Determine the value of x : (x,5) (-4,3) slope=-1????

geerky42 (geerky42):

You can use slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\). Plug in these value and solve for x.

OpenStudy (anonymous):

ok so 3-5/-4-x

geerky42 (geerky42):

yeah, now solve for x

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ x}}\quad ,&{\color{blue}{ 5}})\quad &({\color{red}{ -4}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\implies \cfrac{{\color{blue}{ 3}}-{\color{blue}{ 5}}}{{\color{red}{ -4}}-{\color{red}{ x}}}=1\) solve for "x"

OpenStudy (jdoe0001):

hmmm actually should be -1 so \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ x}}\quad ,&{\color{blue}{ 5}})\quad &({\color{red}{ -4}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\implies \cfrac{{\color{blue}{ 3}}-{\color{blue}{ 5}}}{{\color{red}{ -4}}-{\color{red}{ x}}}=-1\)